OpenStudy (anonymous):

an antique map was found in the attic of a local courthouse. it shows some measurements from a local farm that was divided into 5 parts. some of the measurements have faded with age, so you must find the remaining measurements, as well as calculate the total area and perimeter of the outside property. all lengths are measured in miles.

OpenStudy (phi):

we need more info. can you make a screen shot of the entire question?

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

the question is on the link but its all the way at the bottom znd it has help on it but i dont understand it

OpenStudy (phi):

that is a link to 43 pages.

OpenStudy (anonymous):

you have to scroll all the way down to the bottom

OpenStudy (phi):

which page?

OpenStudy (anonymous):

the very last one

OpenStudy (anonymous):

43

OpenStudy (phi):

ok. It sounds like they want all the angles and sides for each of the triangles

OpenStudy (anonymous):

yeah i think so but im not sure how

OpenStudy (phi):

Mostly the Law of Cosines and Law of Sines Let's start in the lower right. that is a right triangle with sides 3 and 4 use pythagoras to find the hypotenuse.

OpenStudy (anonymous):

so i would do a^2+b^2=c^2

OpenStudy (phi):

yes, and we are looking for c

OpenStudy (anonymous):

4^2+3^2=c^2

OpenStudy (anonymous):

which equals 25

OpenStudy (phi):

c^2 = 25 c = sqr(25)

OpenStudy (anonymous):

5

OpenStudy (phi):

yes, so label side AC with length 5

OpenStudy (anonymous):

ok got it

OpenStudy (phi):

the few pages in front of the last page seem to go through the solution

OpenStudy (anonymous):

i saw that to but i dont really understand how they did that

OpenStudy (phi):

Say the job is to find the sides of all 5 triangles. so far you know the sides of triangle ABC

OpenStudy (anonymous):

ok

OpenStudy (phi):

now look at triangle ACD. we have two sides and we need to find the 3rd side. Do we have enough info to do that ?

OpenStudy (anonymous):

we use law of cosine?

OpenStudy (phi):

yes. If we see side-angle-side with numbers, then we can use the Law of Cosines to find the unknown side. Do you know how to do that to find side AD ?

OpenStudy (anonymous):

\[5^2+3.1623^2-2(5)(3.1623)\cos(71.5651\]

OpenStudy (phi):

that is c^2 what do you get for c^2 and then c ?

OpenStudy (anonymous):

dont you get 35.0-31.623*0.31 and then put that in the calculator?

OpenStudy (anonymous):

and i got 25.1 but i have to square root it

OpenStudy (phi):

I would use the calculator to do the whole computation \[ 5^2+3.1623^2-2(5)(3.1623)\cos(71.565)\]

OpenStudy (phi):

you don't want to round any numbers until the very end.

OpenStudy (anonymous):

i got 25 still

OpenStudy (phi):

so what is side AD ?

OpenStudy (anonymous):

well when i square root 25 i got 5

OpenStudy (phi):

so label side AD 5 it looks like triangle ACD is isosceles (2 sides the same length) that means we know the base angles are both 71.5651 and the angle CAD is 180 - 2*71.5651 I am not sure we we to write that down, but we do know it.

OpenStudy (anonymous):

so angle a is the same as angle c (71.5652)

OpenStudy (phi):

If by angle a and angle c you mean < CDA and < ACD

OpenStudy (anonymous):

yes

OpenStudy (phi):

now triangle ADE we have 2 sides, and an angle, but it is the wrong angle. However, we can find the "middle angle". Any ideas how?

OpenStudy (anonymous):

no i dont know anything about the middle angle

OpenStudy (phi):

Do you see we want to find side DE ? to do that we can use side AE, angle DAE, side AD and the Law of Cosines but we need angle DAE

OpenStudy (phi):

When I see we have sides and angles, and need an angle, I would try using the Law of Sines. What angle can we find if we use the Law of Sines ?

OpenStudy (anonymous):

\[cosc=a^2+b^2-c^2/2ab\]

OpenStudy (phi):

we can't use that because we don't know two things (angle C and side c) I would use the Law of Sines to find angle ADE

OpenStudy (anonymous):

so then sin=a/sinA=b/sinB

OpenStudy (phi):

yes. can you find angle ADE ?

OpenStudy (anonymous):

remind me again is the a the side and the A is the degree right

OpenStudy (phi):

yes, you can only take sines of angles. you can write the Law of Sines as \[ \frac{a}{\sin A} = \frac{b}{\sin B}\] or as \[ \frac{\sin A}{a} = \frac{\sin B}{b}\]

OpenStudy (anonymous):

ok \[\sin=\frac{ 7.8102 }{ sinA }=\frac{ 5 }{ \sin38.845 }\]

OpenStudy (phi):

ok except I would not have that first sin = part just \[ \frac{ 7.8102 }{ \sin A }=\frac{ 5 }{ \sin38.845 }\]

OpenStudy (phi):

typo: \[ \frac{ 7.8102 }{ \sin A }=\frac{ 5 }{ \sin38.8845 } \]

OpenStudy (anonymous):

oh ok and then sin(38.845 is equal to 0.62 so you have \[\frac{ 7.8102 }{ ?sinA}=\frac{ 5 }{ 0.62 }\]

OpenStudy (phi):

it's sin 38.8845 (not 38.845) when you do the calculation you should keep at least 5 decimals (I would keep all that the calculator uses)

OpenStudy (anonymous):

0.6277524996

OpenStudy (phi):

The way I would do it is write it as \[ \frac{\sin A}{7.8102} = \frac{\sin 38.8845}{5} \] and then multiply both sides by 7.8102 to get \[ \sin A =\frac{\sin 38.8845}{5}\cdot 7.8102\] now use a calculator to find sin A

OpenStudy (anonymous):

i put that in and it says error

OpenStudy (phi):

you can use google. type into the google search window sin 38.8845 deg * 7.8102/5=

OpenStudy (anonymous):

okay thank you for everything i got the rest your the best